### v2

# Nonreciprocity of hydrodynamic electron transport in noncentrosymmetric conductors

Abstract
We show that the nonreciprocity of hydrodynamic electron transport in noncentrosymmetric conductors with broken time-reversal symmetry (TRS) is significantly enhanced compared to the disorder-dominated regime. This enhancement is caused by the linear dependence of the viscosity of the electron liquid on the flow velocity, which is allowed in the absence of TRS and Galilean invariance. The resulting nonlinear flows break dynamical similarity and must be characterized by two dimensionless parameters: the Reynolds number and the emergent nonreciprocity number. The latter is linear in velocity but independent of system size. We determine the nonlinear conductance of a Hall bar and show that the nonreciprocal correction to the current can be of comparable magnitude to its reciprocal counterpart.
According to the Onsager reciprocity principle Onsager (1931a); *Onsager-II, the linear two-terminal conductance $G_{0}$ must be invariant under the time-reversal symmetry (TRS), which changes the sign of the magnetic field $\bm{B}$ and the magnetization of the system $\bm{M}$. In contrast, the nonlinear two-terminal transport need not be reciprocal. In noncentrosymmetric conductors, i.e. nonpolar systems lacking inversion symmetry, due to the existence of an invariant that is linear in both the electric field $\bm{E}$ and the TRS breaking pseudovectors $\bm{B}$ or $\bm{M}$, the nonreciprocal contribution to the current density appears already in second order in the electric field $\bm{E}$ Sturman and Fridkin (1992); Rikken et al. (2001); Ivchenko and Spivak (2002). Thus, in a two-terminal setup, the electric current $I$ through the system may be expressed as
$I=G_{0}V+G_{2}V^{2},$ (1)
where $V$ is the voltage bias, and the nonlinear part of the conductance, $G_{2}$, has a nonreciprocal (odd in the TRS-breaking perturbation) part. Nonreciprocity of nonlinear electron transport was the subject of extensive research in the context of mesoscopic systems in the disorder-dominated Drude regime Spivak and Zyuzin (2004); Sánchez and Büttiker (2004); Polianski and Büttiker (2006); De Martino et al. (2006); Deyo et al. (2006), with experiments on quantum dots, carbon nanotubes, and quantum wires Löfgren et al. (2004); Wei et al. (2005); Leturcq et al. (2006); Zumbühl et al. (2006); Angers et al. (2007). In superconducting systems with broken TRS, nonreciprocity arises in both equilibrium current, i.e. the so called superconducting diode effect Nadeem et al. (2023), and the dissipative regime Liu et al. (2024). For a recent review of nonreciprocal transport and optical phenomena in quantum materials see Ref. Nagaosa and Yanase, 2024.
In this article, we develop the theory of nonreciprocal electron transport in noncentrosymmetric conductors with broken TRS in the hydrodynamic regime governed by momentum-conserving electron-electron scattering. The nondissipative effects of TRS breaking in the hydrodynamics of liquids, such as ferrofluids Shliomis (1967); Odenbach (2002), can be described by introducing nondissipative kinetic coefficients of the liquid Lifshitz and Pitaevskii (1981), which are odd under TRS as required by the Onsager symmetry. For instance, significant attention has been devoted to the study of odd, or Hall, viscosity of quantum liquids  Avron (1998); Haldane (2009); Bradlyn et al. (2012); Hoyos and Son (2012); Messica et al. (2025) and active matter systems Banerjee et al. (2017); Aggarwal et al. (2023); Kirkinis and Olvera de la Cruz (2023, 2024) with broken TRS. We consider the effects of TRS breaking on the dissipative nonlinear hydrodynamic transport. In most high-mobility semiconductor heterostructures and graphene devices where the hydrodynamic electron transport has been demonstrated (see, e.g., reviews Narozhny (2022); Fritz and Scaffidi (2024) and references therein), the electron liquid lacks Galilean invariance. Recent experiments indicate that the electron liquid may spontaneously break TRS and inversion symmetry Zhou et al. (2021); Zhang et al. (2024); Lin et al. (2023); Chichinadze et al. (2024).
We find that nonlinear flows of nonreciprocal and conventional liquids are qualitatively different in the following respect. Nonlinear hydrodynamic flows of conventional Newtonian liquids obey dynamical similarity Landau and Lifshitz (1987); Batchelor (1967), which enables modeling of large-scale hydrodynamic phenomena in a lab; flows in homothetic systems can be mapped to one another by rescaling of the hydrodynamic velocity. All equivalent flows have the same Reynolds number $\mathcal{R}=U_{0}L/\nu$, where $U_{0}$ and $L$ are characteristic velocity and length scales, respectively, and $\nu$ is the kinematic viscosity. We show that nonlinear flows of noncentrosymmetric electron liquids with broken time-reversal invariance do not possess dynamical similarity. Besides the Reynolds number, they must be characterized by an additional dimensionless parameter, which is proportional to $U_{0}$ and is odd under time-reversal.
The existence of the nonreciprocal component of $G_{2}$ in Eq. (1) requires breaking not only TRS but also a spatial symmetry, which distinguishes the flows in the forward and reverse directions. Similar to the situation in mesoscopic systems, this symmetry breaking may be associated with the device geometry Andreev and Glazman (2006), e.g. a flow in a funnel. Here we focus on the two-dimensional (2D) Poiseuille flow in the Hall, which does not break the symmetry between the forward and reverse bias. In this case, the required spatial symmetry breaking arises in the noncentrosymmetric electron liquid itself. Below we study nonreciprocal transport for two different types of spatial symmetry breaking relevant to experiments: i) symmetry breaking of a vector type realized in 2D electron liquids with spin-orbit interaction in the presence of an in-plane Zeeman field, ii) symmetry breaking of a tensor type realized in valley-polarized graphene Zhou et al. (2021); Zhang et al. (2024); Lin et al. (2023); Chichinadze et al. (2024). The breaking of forward/reverse symmetry is described in i) by a vector, which is linear in the Zeeman field, and in ii) by an invariant rank-3 tensor that is allowed by the $C_{3}$ rotation symmetry in graphene. In the latter situation, the magnitude of the nonreciprocal conductance depends on the orientation of the Hall bar relative to the graphene lattice.
We show that nonreciprocity of dissipative transport is strongly enhanced in comparison to that in the Drude regime. The enhancement is caused by the linear dependence of the viscosity of the electron liquid on the hydrodynamic velocity $\bm{v}$. Such a dependence is allowed by symmetry in noncentrosymmetric conductors with broken TRS. Its existence can be established in the framework of the Boltzmann equation for model systems of vector and tensor type symmetry breaking mentioned above. In a general situation, the liquid viscosity is described by a rank-4 tensor $\hat{\eta}$, which relates the viscous tress tensor
$\sigma_{ij}=\eta_{ijkl}V_{kl},\quad V_{ij}=\frac{1}{2}\left(\partial_{j}v_{i}+\partial_{i}v_{j}\right),$ (2)
to the strain rate tensor $V_{ij}$. For Galilean-invariant liquids, the viscosity tensor cannot depend on the flow velocity, and for isotropic incompressible liquids in two dimensions it can be reduced to the shear, $\eta$, and odd, $\eta_{H}$, viscosities Avron (1998). When both TRS and Galilean invariance are broken, a linear in the flow velocity correction to the viscous stress tensor is permitted,
$\delta\sigma_{ij}=$ $\,T_{ijklm}v_{k}\partial_{l}v_{m}.$ (3)
In vector-type symmetry breaking, where the TRS is broken by an external magnetic field $\bm{B}$, the rank-5 tensor $T_{ijklm}$ may be constructed using $\bm{B}$, and the invariant tensors of Levi-Civita, $\epsilon_{ij}$, and Kronecker, $\delta_{ij}$. This yields
$\delta\sigma_{ij}=$ $\eta\left[V_{ij}\left(\alpha\delta_{kl}B_{l}v_{k}+\widetilde{\alpha}\epsilon_{lk}B_{k}v_{l}\right)\right.$
$\left.+\beta v_{i}B_{k}\partial_{k}v_{j}+\widetilde{\beta}B_{i}v_{k}\partial_{k}v_{j}+\ldots\right].$ (4)
Here we retained only several representative terms in the constitutive law and introduced phenomenological parameters $\alpha,\beta,\ldots$ whose values must be determined from a microscopic theory.
In tensor-type symmetry breaking, which is realized in the recently discovered quarter-metal state in rhombohedral trilayer graphene Zhou et al. (2021), the TRS breaking is caused by valley and spin polarization. Consequently, the system does not possess an in-plane vector breaking the TRS. However, due to the trigonal warping of the electron spectrum, the rotational symmetry of the electron liquid is lowered to $C_{3}$, which allows for an invariant rank-3 tensor $R_{ijk}$. The linear in $\bm{v}$ tensor $T_{ijklm}$ in Eq. (3) can be constructed using $R_{ijk}$, $\delta_{ij}$, and $\epsilon_{ij}$. It is more convenient to express this tensor relation in the chiral complex coordinates $z=x+iy$ and $\bar{z}=x-iy$. In these coordinates, a traceless stress tensor of an incompressible liquid is described by a single element $\delta\sigma_{zz}=\delta\sigma_{xx}-\delta\sigma_{yy}+2i\delta\sigma_{xy}$, and the invariant rank-3 tensor $R$ has only two non-vanishing components: $r\equiv R_{zzz}$ and $\bar{r}\equiv R_{\bar{z}\bar{z}\bar{z}}$ Landau and Lifshitz (1986) (since in Cartesian coordinates the tensor $R$ is real, $\bar{r}$ is the complex-conjugate of $r$). Upon a rotation by $32\pi/$ 3 the chiral coordinates are transformed as $z\to\epsilon z$, $\bar{z}\to\epsilon^{-1}\bar{z}$, where $\epsilon=e^{2\pi i/3}$. Therefore, in the complex coordinates the nonreciprocal part of the viscous stress tensor in Eq. (Nonreciprocity of hydrodynamic electron transport in noncentrosymmetric conductors) has the form
$\delta\sigma_{zz}=$ $\,\eta rV_{\bar{z}\bar{z}}v_{z},\quad\delta\sigma_{\bar{z}\bar{z}}=\,\eta\bar{r}V_{zz}v_{\bar{z}}.$ (5)
The characteristic ratio of the nonreciprocal stress in Eqs. (Nonreciprocity of hydrodynamic electron transport in noncentrosymmetric conductors), and Eq. (5) to the standard viscous stress $(\sigma_{0})_{ij}=2\eta V_{ij}$ in a given flow defines a dimensionless parameter, which is proportional to the characteristic flow velocity $U_{0}$ but is distinct from the Reynolds number,
$\mathcal{N}\sim\frac{\delta\hat{\sigma}}{\hat{\sigma}_{0}}\propto U_{0}.$ (6)
We refer to $\mathcal{N}$ as the nonreciprocity number. It can be interpreted as the ratio of the typical flow velocity $U_{0}$ to the characteristic velocity scale associated with TRS breaking. In vector-type symmetry breaking the latter is proportional to the strength of spin-orbit coupling and the Zeeman field. In tensor-type, assuming that the degree of valley polarization is of order unity, the characteristic velocity is given by the Fermi velocity modulation caused by the trigonal warping. Since $\mathcal{N}$ is proportional to the flow velocity, but independent of system size, it breaks dynamical similarity of nonlinear flows.
Figure 1: 2D nonreciprocal flow in a Hall bar of width $2d$. Nonreciprocal velocity profile $\bm{v}=u(y)\hat{\mathbf{x}}$ is controlled by the nonreciprocity number $\mathcal{N}$ in (6), which depends on the orientation of the in-plane magnetic field $B$ in vector-type symmetry breaking, and on the valley polarization and the orientation of the graphene lattice in tensor-type. Panel (a) displays the velocity profile (10) for $\mathcal{N}=0.3$ and $0.9$, while panel (b) displays the velocity profiles for the time-reversed values $\mathcal{N}=-0.3$ and $-0.9$. In both cases the Poiseuille profile is restored for $\mathcal{N}\rightarrow 0$.
To obtain quantitative results for nonreciprocal electron transport in the hydrodynamic regime, we consider the experimentally relevant two-dimensional Poiseuille flow in the Hall bar geometry shown in Fig. 1. Transport of momentum in a steady-state hydrodynamic flow is described by the force balance equation Landau and Lifshitz (1987)
$neE_{i}=\partial_{j}\Pi_{ij},$ (7)
where $n$ is the density of electrons, $e$ is the electron charge, $\bm{E}$ is the electric field, and $\Pi_{ij}$ is the momentum flux tensor density. The latter is conventionally expressed in terms of the pressure $P\delta_{ij}$ and the viscous stress tensor $\sigma_{ij}$, as
$\Pi_{ij}=P\delta_{ij}+\rho v_{i}v_{j}-2\eta V_{ij}-\delta\sigma_{ij}.$ (8)
In the absence of Galilean invariance, the expression $P\delta_{ij}+\rho v_{i}v_{j}$ for the equilibrium part of the momentum flux tensor in Eq. (8) should be viewed as a formal expansion of momentum flux density in an equilibrium state of uniform flow to second order in the powers of the flow velocity $\bm{v}$ with higher-order terms omitted.
For a current flowing through a channel of width $2d$, as displayed in Fig. 1, the velocity field has only one nonzero component $\bm{v}=u(y)\hat{\mathbf{x}}$, which depends only on the transverse coordinate $y$. The viscous force density is fully described by the single component of the viscous stress, $\sigma_{xy}(y)$. As a result, Eq. (7) reduces to
$\frac{en\mathcal{E}}{U_{0}}+\partial_{y}\left[\eta\left(1+\mathcal{N}\frac{u}{U_{0}}\right)\partial_{y}\frac{u}{U_{0}}\right]=0,$ (9)
where we absorbed the pressure gradient $\partial_{x}P$ into the electromotive force (EMF) density $ne\mathcal{E}=neE-\partial_{x}P$. The latter is related to the voltage drop, $\mathcal{E}=V/L$, with $L$ being the length of the Hall bar  [^1] It is worth noting that for Galilean-invariant liquids, the Navier-Stokes equation predicts a linear current-voltage dependence for the Poiseuille flow at arbitrary Reynolds numbers. Therefore, in this case all nonlinearity is nonreciprocal.. For vector- and tensor- type of symmetry breaking the value of the nonreciprocity parameter $\mathcal{N}$ in Eq. (9) is easily determined from Eqs. (Nonreciprocity of hydrodynamic electron transport in noncentrosymmetric conductors) and (5). For vector-type the precise form of $\mathcal{N}$ depends on the mutual orientation of the field with respect to the flow velocity. For tensor-type $\mathcal{N}=(r+\bar{r})U_{0}/2$. Notice that a rotation of crystalline axes by an angle $\theta$ relative to the flow changes $r\to re^{3i\theta}$. Therefore, the combination $r+\bar{r}$ in the nonreciprocity parameter exhibits a periodic modulation $\propto\cos(3[\theta-\theta_{0}])$. This is consistent with the three-fold rotation symmetry observed in nonreciprocity measurements in valley-polarized multilayer graphene Lin et al. (2023); Chichinadze et al. (2024).
Equation (9) can be readily integrated with the integration constants fixed by the boundary condition imposed on the flow. For simplicity, we apply no-slip conditions $u(\pm d)=0$. The more general case of a flow with a finite slip length does not alter the essential physics underlying nonreciprocity. Thus, the profile of the flow can be found in the form
$u(y)=\frac{U_{0}}{\mathcal{N}}\left[\sqrt{1+\mathcal{N}\left(1-\frac{y^{2}}{d^{2}}\right)}-1\right],$ (10)
where we introduced $U_{0}=en\mathcal{E}d^{2}/\eta$. Equation (10) is valid for both positive and negative values of $\mathcal{N}$. At $\mathcal{N}\to 0$ it reproduces the Poiseuille flow profile, see Fig. 1 for illustration.
From the flow profile, we can compute the total current  [^2] In systems lacking Galilean invariance, the total current density consists of the usual convective term, $enu$, and an additional contribution arising from the finite intrinsic conductivity Li et al. (2022). Assuming the particle density is sufficiently high, we neglect the latter.. The resulting expression for the total current ($B>0$) takes a relatively simple form
[…]
where the dimensionless function is given by
$f(\mathcal{N})=\frac{3}{2\mathcal{N}^{\frac{3}{2}}}\left[(1+\mathcal{N})\arcsin\sqrt{\frac{\mathcal{N}}{1+\mathcal{N}}}-\sqrt{\mathcal{N}}\right].$ (12)
The field-independent linear conductance, $G_{0}=(I/V)_{\mathcal{N}\to 0}$, takes the value
$G_{0}=\frac{2e^{2}}{3}(nd^{2})\frac{n}{\eta}\frac{d}{L}.$ (13)
It is inversely proportional to the viscosity of the electron fluid, which is the manifestation of the Gurzhi effect Gurzhi (1968). This behavior was confirmed experimentally Krishna Kumar et al. (2017); Krebs et al. (2023); Zeng et al. (2024). Working to leading order in the strength of the TRS breaking perturbation, $\mathcal{N}\ll 1$, we extract the nonlinear nonreciprocal correction to conductance of the system. In the notations of Eq. (1) it takes the form
$G_{2}=-\frac{G_{0}}{5}\frac{nd^{2}}{\eta L}\left\{\begin{array}[]{cc}(e\alpha B),&\text{vector-type},\\ e|r|\cos(3[\theta-\theta_{0}]),&\text{tensor-type}.\end{array}\right.$ (14)
This is obtained by expanding the expression $I/I_{0}=f(\mathcal{N})\approx 1-\mathcal{N}/5$ in (11), where $f$ is given by Eq. (12), and keeping only leading order terms in the nonreciprocity number $\mathcal{N}$. For simplicity, in vector-type symmetry breaking, we also took a field orientation $\bm{B}=B\hat{\mathbf{x}}\parallel\bm{v}$ so that nonreciprocity number simplifies to $\mathcal{N}=\alpha BU_{0}$. In Fig. 2 we display the function $f(\mathcal{N})$ vs. $\mathcal{N}$, along with the limiting cases of weak and strong nonreciprocity.
Figure 2: Plot of the dimensionless current $I/I_{0}=f(\mathcal{N})$ in Eq. (11) (continuous curve) where $f$ is given by Eq. (12), vs. the nonreciprocity number for $\mathcal{N}>0$, cf. Fig. 1(a). The dashed curve on the left denotes its linear approximation, $I/I_{0}\approx 1-\mathcal{N}/5$, whose nonreciprocal correction to the conductance was calculated in Eq. (14). The dashed curve on the right denotes its high $\mathcal{N}$ asymptotic expansion $I/I_{0}\approx 3\pi/(4\sqrt{\mathcal{N}})$. Inset: Plot of the dimensionless current $I/I_{0}$ calculated for the negative branch of the flow profile displayed in Fig. 1(b).
We now compare the magnitude of nonreciprocity in the hydrodynamic regime to that in the disorder-dominated regime. In the latter case, the ratio of the nonreciprocal to reciprocal current can be expressed as
$\frac{\delta I_{\text{dis}}}{I_{0}}\sim\varsigma\frac{e\mathcal{E}\ell_{\text{ei}}}{\epsilon_{\text{F}}},$ (15)
where $\ell_{\text{ei}}=v_{\text{F}}\tau_{\text{ei}}$ is the disorder-limited electron mean-free path, and $\varsigma$ is a dimensionless strength of TRS breaking. For example, for carbon nanotubes considered in Ref. Spivak and Zyuzin, 2004, $\varsigma\sim\Phi/\Phi_{0}$, where $\Phi$ is the magnetic flux through the nanotube and $\Phi_{0}=hc/e$ is the flux quantum. In the hydrodynamic regime, the corresponding estimate, for the relative magnitude of the nonreciprocal current is significantly enhanced in comparison to its disorder-dominated counterpart. In valley-polarized graphene, the degree of nonreciprocity of the electron spectrum depends on the energy scale $\Delta$ associated with the trigonal warping. The corresponding dimensionless parameter $\varsigma\sim\Delta/\epsilon_{\text{F}}$, may in principle reach values of order unity. For the case of 2D electron liquid with Rashba spin-orbit coupling, this parameter can be estimated as a ratio of the Zeeman energy to the spin-orbit splitting. Using the estimate for the typical flow velocity from Eq. (10), $U_{0}\sim\frac{d^{2}}{\eta}en\mathcal{E}$, and estimating the shear viscosity in the Fermi liquid regime as $\eta\sim np_{\text{F}}\ell_{\text{ee}}$, where $\ell_{\text{ee}}=v_{\text{F}}\tau_{\text{ee}}$ is the inelastic mean free path, we obtain
$\frac{\delta I_{\text{hydro}}}{I_{0}}\sim\varsigma\frac{d}{\ell_{\text{ee}}}\frac{e\mathcal{E}d}{\epsilon_{\text{F}}}.$ (16)
Comparing Eqs. (16) and (15) we see that in the hydrodynamic regime the elastic electron mean free path is effectively replaced by the channel width, $\ell_{\text{ei}}\to d$, and an additional large factor $d/\ell_{\text{ee}}\gg 1$ arises from the dependence of the flow velocity on the channel width. Thus, the relative nonreciprocal correction in the hydrodynamic regime is enhanced in comparison to the disorder-dominated regime by a large factor
$\frac{\delta I_{\text{hydro}}}{\delta I_{\text{dis}}}\sim\frac{d^{2}}{\ell_{\mathrm{ee}}\ell_{\mathrm{ei}}}.$ (17)
In the presence of impurity scattering in the bulk, the hydrodynamic approach used above is valid provided $d$ does not exceed the Gurzhi length, $\ell_{\text{G}}=\sqrt{\ell_{\text{ei}}\ell_{\text{ee}}}$. For wide channels ($d>\ell_{\text{G}}$), electron flow is dominated by the scattering in the bulk of the device, leading to a transition from hydrodynamic Poiseuille flow to Ohmic behavior. As expected, at the crossover boundary between the two regimes, $d\sim\ell_{\text{G}}$, the results for nonreciprocity in Eqs. (16) and (15) match.
We thank B. Spivak for useful discussions. The work of L. B. and A. V. A. was supported by the National Science Foundation Grant No. DMR-2424364, the Thouless Institute for Quantum Matter (TIQM), and the College of Arts and Sciences at the University of Washington. The work of A. L. was supported by NSF Grant No. DMR-2203411 and H. I. Romnes Faculty Fellowship provided by the University of Wisconsin-Madison Office of the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation.

### v3

@@
 This enhancement is caused by the linear dependence of the viscosity of the electron liquid on the flow velocity, which is allowed in the absence of TRS and Galilean invariance.
-The resulting nonlinear flows break dynamical similarity and must be characterized by two dimensionless parameters: the Reynolds number and the emergent nonreciprocity number.
+The resulting nonlinear flows break dynamical similarity and thus, in addition to the Reynolds number, they are characterized by a second dimensionless parameter, the nonreciprocity number.
 The latter is linear in velocity but independent of system size.
@@
 Nagaosa and Yanase, 2024.
-In this article, we develop the theory of nonreciprocal electron transport in noncentrosymmetric conductors with broken TRS in the hydrodynamic regime governed by momentum-conserving electron-electron scattering.
+In this Letter, we develop the theory of nonreciprocal electron transport in noncentrosymmetric conductors with broken TRS in the hydrodynamic regime governed by momentum-conserving electron-electron scattering.
 The nondissipative effects of TRS breaking in the hydrodynamics of liquids, such as ferrofluids Shliomis (1967); Odenbach (2002), can be described by introducing nondissipative kinetic coefficients of the liquid Lifshitz and Pitaevskii (1981), which are odd under TRS as required by the Onsager symmetry.
@@
 (2024).
-We find that nonlinear flows of nonreciprocal and conventional liquids are qualitatively different in the following respect.
+We establish that nonlinear flows of nonreciprocal liquids differ qualitatively from their conventional counterparts in the following respect.
 Nonlinear hydrodynamic flows of conventional Newtonian liquids obey dynamical similarity Landau and Lifshitz (1987); Batchelor (1967), which enables modeling of large-scale hydrodynamic phenomena in a lab; flows in homothetic systems can be mapped to one another by rescaling of the hydrodynamic velocity.
 All equivalent flows have the same Reynolds number $\mathcal{R}=U_{0}L/\nu$, where $U_{0}$ and $L$ are characteristic velocity and length scales, respectively, and $\nu$ is the kinematic viscosity.
-We show that nonlinear flows of noncentrosymmetric electron liquids with broken time-reversal invariance do not possess dynamical similarity.
+In contrast, we show that nonlinear flows of noncentrosymmetric electron liquids with broken time-reversal invariance do not possess dynamical similarity.
 Besides the Reynolds number, they must be characterized by an additional dimensionless parameter, which is proportional to $U_{0}$ and is odd under time-reversal.
@@
 a flow in a funnel.
-Here we focus on the two-dimensional (2D) Poiseuille flow in the Hall, which does not break the symmetry between the forward and reverse bias.
+Here we focus on the two-dimensional (2D) Poiseuille flow in the Hall bar, which does not break the symmetry between the forward and reverse bias.
 In this case, the required spatial symmetry breaking arises in the noncentrosymmetric electron liquid itself.
@@
 Its existence can be established in the framework of the Boltzmann equation for model systems of vector and tensor type symmetry breaking mentioned above.
-In a general situation, the liquid viscosity is described by a rank-4 tensor $\hat{\eta}$, which relates the viscous tress tensor
+In a general situation, the liquid viscosity is described by a rank-4 tensor $\hat{\eta}$, which relates the viscous stress tensor
 $\sigma_{ij}=\eta_{ijkl}V_{kl},\quad V_{ij}=\frac{1}{2}\left(\partial_{j}v_{i}+\partial_{i}v_{j}\right),$ (2)
@@
 $\left.+\beta v_{i}B_{k}\partial_{k}v_{j}+\widetilde{\beta}B_{i}v_{k}\partial_{k}v_{j}+\ldots\right].$ (4)
-Here we retained only several representative terms in the constitutive law and introduced phenomenological parameters $\alpha,\beta,\ldots$ whose values must be determined from a microscopic theory.
+Here we retained only several representative terms in the constitutive law and introduced phenomenological parameters $\alpha,\beta,\ldots$ whose values must be determined from a microscopic theory  [^1] In the framework of the Boltzmann transport equation, these parameters can be obtained following the method of Enskog (see e.g.
+§6 of Ref.
+Lifshitz and Pitaevskii (1981)).
+It is important that for non-parabolic electron dispersion $\epsilon(\bm{p})$, the locally-equilibrium distribution function in the presence of the hydrodynamic velocity $\bm{v}$, $n_{0}(\bm{p})=\left[e^{(\epsilon(\bm{p})-\bm{p}\cdot\bm{v}-\mu)/T}+1\right]^{-1}$, does not amount to a shift of momentum and the chemical potential $\mu$, but alters the shape of the Fermi surface.
+The distortion of the Fermi surface caused by $\bm{v}$ modifies all the kinetic coefficients of the electron liquid, including the viscosity.
+In electronic liquids with broken inversion and time-reversal symmetry, the modification of viscosity starts from the linear order in $\bm{v}$..
 In tensor-type symmetry breaking, which is realized in the recently discovered quarter-metal state in rhombohedral trilayer graphene Zhou et al.
@@
 In these coordinates, a traceless stress tensor of an incompressible liquid is described by a single element $\delta\sigma_{zz}=\delta\sigma_{xx}-\delta\sigma_{yy}+2i\delta\sigma_{xy}$, and the invariant rank-3 tensor $R$ has only two non-vanishing components: $r\equiv R_{zzz}$ and $\bar{r}\equiv R_{\bar{z}\bar{z}\bar{z}}$ Landau and Lifshitz (1986) (since in Cartesian coordinates the tensor $R$ is real, $\bar{r}$ is the complex-conjugate of $r$).
-Upon a rotation by $32\pi/$ 3 the chiral coordinates are transformed as $z\to\epsilon z$, $\bar{z}\to\epsilon^{-1}\bar{z}$, where $\epsilon=e^{2\pi i/3}$.
-Therefore, in the complex coordinates the nonreciprocal part of the viscous stress tensor in Eq.
-(Nonreciprocity of hydrodynamic electron transport in noncentrosymmetric conductors) has the form
-$\delta\sigma_{zz}=$ $\,\eta rV_{\bar{z}\bar{z}}v_{z},\quad\delta\sigma_{\bar{z}\bar{z}}=\,\eta\bar{r}V_{zz}v_{\bar{z}}.$ (5)
-The characteristic ratio of the nonreciprocal stress in Eqs.
-(Nonreciprocity of hydrodynamic electron transport in noncentrosymmetric conductors), and Eq.
-(5) to the standard viscous stress $(\sigma_{0})_{ij}=2\eta V_{ij}$ in a given flow defines a dimensionless parameter, which is proportional to the characteristic flow velocity $U_{0}$ but is distinct from the Reynolds number,
-$\mathcal{N}\sim\frac{\delta\hat{\sigma}}{\hat{\sigma}_{0}}\propto U_{0}.$ (6)
+Upon a rotation by $2\pi/3$ the chiral coordinates are transformed as $z\to\epsilon z$, $\bar{z}\to\epsilon^{-1}\bar{z}$, where $\epsilon=e^{2\pi i/3}$.
+Therefore, we get
+$\delta\sigma_{zz}=$ $\,\eta rV_{\bar{z}\bar{z}}v_{z},\quad\delta\sigma_{\bar{z}\bar{z}}=\,\eta\bar{r}V_{zz}v_{\bar{z}},$ (5)
+for the nonvanishing components of the nonreciprocal viscous stress tensor in Eq. (Nonreciprocity of hydrodynamic electron transport in noncentrosymmetric conductors).
+For a given flow, the ratio of the characteristic nonreciprocal stress in Eqs.
+(Nonreciprocity of hydrodynamic electron transport in noncentrosymmetric conductors) and (5) to the standard viscous stress $(\sigma_{0})_{ij}=2\eta V_{ij}$ defines a dimensionless parameter
+$\mathcal{N}\sim\frac{\delta\sigma}{\sigma_{0}}\propto U_{0},$ (6)
+proportional to the characteristic flow velocity $U_{0}$ but distinct from the Reynolds number.
 We refer to $\mathcal{N}$ as the nonreciprocity number.
@@
 Nonreciprocal velocity profile $\bm{v}=u(y)\hat{\mathbf{x}}$ is controlled by the nonreciprocity number $\mathcal{N}$ in (6), which depends on the orientation of the in-plane magnetic field $B$ in vector-type symmetry breaking, and on the valley polarization and the orientation of the graphene lattice in tensor-type.
-Panel (a) displays the velocity profile (10) for $\mathcal{N}=0.3$ and $0.9$, while panel (b) displays the velocity profiles for the time-reversed values $\mathcal{N}=-0.3$ and $-0.9$.
+Panel (a) displays the velocity profile (10) for $\mathcal{N}=0.3$ and $0.9$, while panel (b) displays the velocity profiles for the time-reversed values $\mathcal{N}=-0.3$ and 0.9-0.9.
 In both cases the Poiseuille profile is restored for $\mathcal{N}\rightarrow 0$.
@@
 The viscous force density is fully described by the single component of the viscous stress, $\sigma_{xy}(y)$.
-As a result, Eq.
+Its nonreciprocal part $\delta\sigma_{xy}=\eta Nu\partial_{y}u$ is characterized by a single parameter $N$, which has units of inverse velocity.
+For a characteristic flow velocity $U_{0}$, the nonreciprocity number is then defined as $\mathcal{N}=NU_{0}$, and Eq.
 (7) reduces to
@@
 where we absorbed the pressure gradient $\partial_{x}P$ into the electromotive force (EMF) density $ne\mathcal{E}=neE-\partial_{x}P$.
-The latter is related to the voltage drop, $\mathcal{E}=V/L$, with $L$ being the length of the Hall bar  [^1] It is worth noting that for Galilean-invariant liquids, the Navier-Stokes equation predicts a linear current-voltage dependence for the Poiseuille flow at arbitrary Reynolds numbers.
-Therefore, in this case all nonlinearity is nonreciprocal..
+The latter is related to the voltage drop, $\mathcal{E}=V/L$, with $L$ being the length of the Hall bar.
+For Poiseuille flow, the convective nonlinear terms in the Navier–Stokes equation do not affect the velocity profile Landau and Lifshitz (1987).
+Therefore, the nonlinear two-terminal conductance arises only from the velocity dependence of the viscosity, and is entirely nonreciprocal.
 For vector- and tensor- type of symmetry breaking the value of the nonreciprocity parameter $\mathcal{N}$ in Eq.
-(9) is easily determined from Eqs.
+(9) is determined from Eqs.
 (Nonreciprocity of hydrodynamic electron transport in noncentrosymmetric conductors) and (5).
 For vector-type the precise form of $\mathcal{N}$ depends on the mutual orientation of the field with respect to the flow velocity.
-For tensor-type $\mathcal{N}=(r+\bar{r})U_{0}/2$.
+For tensor-type we have $\mathcal{N}=(r+\bar{r})U_{0}/2$.
 Notice that a rotation of crystalline axes by an angle $\theta$ relative to the flow changes $r\to re^{3i\theta}$.
@@
 The more general case of a flow with a finite slip length does not alter the essential physics underlying nonreciprocity.
-Thus, the profile of the flow can be found in the form
-$u(y)=\frac{U_{0}}{\mathcal{N}}\left[\sqrt{1+\mathcal{N}\left(1-\frac{y^{2}}{d^{2}}\right)}-1\right],$ (10)
-where we introduced $U_{0}=en\mathcal{E}d^{2}/\eta$.
-Equation (10) is valid for both positive and negative values of $\mathcal{N}$.
-At $\mathcal{N}\to 0$ it reproduces the Poiseuille flow profile, see Fig.
-1 for illustration.
-From the flow profile, we can compute the total current  [^2] In systems lacking Galilean invariance, the total current density consists of the usual convective term, $enu$, and an additional contribution arising from the finite intrinsic conductivity Li et al.
+Choosing $U_{0}$ as the Poiseuille flow velocity at the center of the channel, $U_{0}=en\mathcal{E}d^{2}/\eta$, we express the profile of the flow in the form
+$u(y)=\frac{U_{0}}{\mathcal{N}}\left[\sqrt{1+\mathcal{N}\left(1-\frac{y^{2}}{d^{2}}\right)}-1\right].$ (10)
+Equation (10) is valid for both positive and negative values of $\mathcal{N}$. At $\mathcal{N}\to 0$ it reproduces the Poiseuille flow profile, see Fig. 1 for illustration.
+From the flow profile, we can compute the total current  [^2] In systems lacking Galilean invariance, the total current density consists of the usual convective term, enuenu, and an additional contribution arising from the finite intrinsic conductivity Li et al.
 (2022).
@@
 The resulting expression for the total current ($B>0$) takes a relatively simple form
-[…]
+$I=\int_{-d}^{+d}\!\!enu(y)dy=I_{0}f(\mathcal{N}),\quad I_{0}=\frac{2}{3}\frac{(en)^{2}\mathcal{E}d^{3}}{\eta},$ (11)
 where the dimensionless function is given by
@@
 (2024).
-Working to leading order in the strength of the TRS breaking perturbation, $\mathcal{N}\ll 1$, we extract the nonlinear nonreciprocal correction to conductance of the system.
+The applicability of Eqs.
+(11) and (12) at increasing bias is limited by the condition of stability of the laminar flow.
+A different condition, $\mathcal{N}\lesssim 1$, arises from expanding the viscous stress tensor to linear order in the velocity.
+Working to leading order in $\mathcal{N}$, we extract the nonlinear nonreciprocal correction to conductance of the system.
 In the notations of Eq.
@@
 (12), and keeping only leading order terms in the nonreciprocity number $\mathcal{N}$.
-For simplicity, in vector-type symmetry breaking, we also took a field orientation $\bm{B}=B\hat{\mathbf{x}}\parallel\bm{v}$ so that nonreciprocity number simplifies to $\mathcal{N}=\alpha BU_{0}$.
+For TRS breaking caused by an in-plane magnetic field $\bm{B}$, the nonreciprocity number is $\mathcal{N}=\alpha BU_{0}$, where $\alpha$ depends on the spin-orbit coupling and the angle between $\bm{B}$ and the flow direction.
 In Fig.
@@
 Figure 2: Plot of the dimensionless current $I/I_{0}=f(\mathcal{N})$ in Eq.
-(11) (continuous curve) where $f$ is given by Eq.
+(11) where $f$ is given by Eq.
 (12), vs.
-the nonreciprocity number for $\mathcal{N}>0$, cf.
+the nonreciprocity number for positive branch ($\mathcal{N}>0$) and negative branch ($\mathcal{N}<0$) of the flow profile, cf.
 Fig.
-1(a).
-The dashed curve on the left denotes its linear approximation, $I/I_{0}\approx 1-\mathcal{N}/5$, whose nonreciprocal correction to the conductance was calculated in Eq.
+1.
+The dashed line denotes its linear approximation, $I/I_{0}\approx 1-\mathcal{N}/5$, whose nonreciprocal correction to the conductance was calculated in Eq.
 (14).
-The dashed curve on the right denotes its high $\mathcal{N}$ asymptotic expansion $I/I_{0}\approx 3\pi/(4\sqrt{\mathcal{N}})$.
-Inset: Plot of the dimensionless current $I/I_{0}$ calculated for the negative branch of the flow profile displayed in Fig.
-1(b).
+At fixed bias, this shows the dependence of the nonlinear conductance on the TRS breaking strength.
 We now compare the magnitude of nonreciprocity in the hydrodynamic regime to that in the disorder-dominated regime.
@@
 For the case of 2D electron liquid with Rashba spin-orbit coupling, this parameter can be estimated as a ratio of the Zeeman energy to the spin-orbit splitting.
-Using the estimate for the typical flow velocity from Eq.
-(10), $U_{0}\sim\frac{d^{2}}{\eta}en\mathcal{E}$, and estimating the shear viscosity in the Fermi liquid regime as $\eta\sim np_{\text{F}}\ell_{\text{ee}}$, where $\ell_{\text{ee}}=v_{\text{F}}\tau_{\text{ee}}$ is the inelastic mean free path, we obtain
+Considering the expression $U_{0}\sim\frac{d^{2}}{\eta}en\mathcal{E}$ for the typical flow velocity and estimating the shear viscosity in the Fermi liquid regime as $\eta\sim np_{\text{F}}\ell_{\text{ee}}$, where $\ell_{\text{ee}}=v_{\text{F}}\tau_{\text{ee}}$ is the inelastic mean free path, we obtain
 $\frac{\delta I_{\text{hydro}}}{I_{0}}\sim\varsigma\frac{d}{\ell_{\text{ee}}}\frac{e\mathcal{E}d}{\epsilon_{\text{F}}}.$ (16)
@@
 was supported by NSF Grant No.
-DMR-2203411 and H.
+DMR-2452658 and H.
 I.
 Romnes Faculty Fellowship provided by the University of Wisconsin-Madison Office of the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation.
+No data were created or analyzed in this study.